OpenStudy (anonymous):

The square root of 8x times the square root of 8x

OpenStudy (nikita2):

\[\sqrt{8x}\sqrt{8x}\] ?

OpenStudy (nikita2):

= 8x and x>=0

OpenStudy (anonymous):

Can you explain better how you got that?

OpenStudy (turingtest):

\[\sqrt a\times\sqrt a=(\sqrt a)^2=a\]

OpenStudy (anonymous):

Actually the original problem is the square root of 5 divided by the square root of 8x. Can you show the whole process because I'm really confused

OpenStudy (anonymous):

But i understand how the square roots cancel out

OpenStudy (nikita2):

\[\sqrt{5}/\sqrt{8x} = \sqrt{\left(\begin{matrix}5 \\ 8x\end{matrix}\right)}\]

OpenStudy (anonymous):

I need to rationalize the denominator

OpenStudy (turingtest):

\[\frac{\sqrt5}{\sqrt{8x}}=\frac{\sqrt5}{\sqrt{2^3x}}\cdot\frac{\sqrt{2x}}{\sqrt{2x}}=\frac{\sqrt{10x}}{4x}\]

OpenStudy (anonymous):

Why don't i multiply the numerator and denominator by the swuare root of 8x?? why do i have to break it up

OpenStudy (turingtest):

\[\frac{\sqrt5}{\sqrt{8x}}=\frac{\sqrt5}{\sqrt{8x}}\cdot\frac{\sqrt{8x}}{\sqrt{8x}}=\frac{\sqrt{40x}}{\sqrt{64x^2}}=\frac{2\sqrt{10x}}{8x}=\frac{\sqrt{10x}}{4x}\]

OpenStudy (anonymous):

wait i thought you said \[\sqrt{8x } * \sqrt{8x} = 8x\]

OpenStudy (turingtest):

yeah it does, look at the denominator. it winds up as 8x, then simplifies to 4x

OpenStudy (anonymous):

and on the last step you just divided the top and bottum by 2 to get rid of it?

OpenStudy (turingtest):

yeah

OpenStudy (anonymous):

ok thank you!

OpenStudy (turingtest):

no prob